# Data-driven rainfall prediction at a regional scale: a case study with Ghana

### v2

@@ Abstract
With a warming planet, tropical regions are expected to experience the brunt of climate change, with more intense and more volatile rainfall events. Currently, state-of-the-art numerical weather prediction (NWP) models are known to struggle to produce skillful rainfall forecasts in tropical regions of Africa. There is thus a pressing need for improved rainfall forecasting in these regions. Over the last decade or so, the increased availability of large-scale meteorological datasets and the development of powerful machine learning models have opened up new opportunities for data-driven weather forecasting. Focusing on Ghana in this study, we use these tools to develop two U-Net convolutional neural network (CNN) models, to predict 24h rainfall at 12h and 30h lead-time. The models were trained using data from the ERA5 reanalysis dataset, and the GPM-IMERG dataset. A special attention was paid to interpretability. We developed a novel statistical methodology that allowed us to probe the relative importance of the meteorological variables input in our model, offering useful insights into the factors that drive precipitation in the Ghana region. Empirically, we found that our 12h lead-time model has performances that match, and in some accounts are better than the 18h lead-time forecasts produced by the ECMWF (as available in the TIGGE dataset). We also found that combining our data-driven model with classical NWP further improves forecast accuracy.
@@ 1 Introduction
The precipitation of liquid water from Earth’s atmosphere to the surface is one of the most crucial atmospheric processes shaping life on earth (Box and Box, 2015). Lack of rainfall leads to drought, and too much rainfall often leads to flooding, two natural disasters that have been plaguing human societies since the dawn of time. The spatio-temporal distribution and intensity of precipitation events also hold significant implications for the smooth running of human societies, from day-to-day events, to large-scale urban and agricultural plans. As such, precipitation prediction is an essential tool for planning and decision making. However, rain and other precipitations are notoriously difficult to predict (Tapiador et al., 2019). Unlike most atmospheric variables, rainfall amounts are spatio-temporally non-smooth, and this poses significant challenge to modeling, with many models often based (implicitly or explicitly) on smoothness assumptions.
@@
The situation is particularly problematic under the tropics where rainfall are dominated by multiscale convective flows that are currently challenging to represent in numerical weather prediction (NWP) models (Becker et al. (2021)). In tropical West Africa, this issue is compounded with the lack of quality measurements. As a consequence state-of-the-art NWP models such as the European Center for Mid-Range Weather Forecasting (ECMWF) models are known to struggle to produce skillful forecasts in these regions (Vogel et al. (2020); Kniffka et al. (2020); Rojas-Campos et al. (2023)). At the same time, regions of sub-Sahara Africa are projected to experience the brunt of climate change, with more intense and more volatile precipitation as Earth’s atmosphere warms (WMO, 2024). For instance, flooding from the 2024 raining season in West and Central Africa have claimed the lives of more than a thousand, and displaced close to a million people [^1] https://www.nytimes.com/2024/09/15/world/africa/floods-africa.html?smid=url-share.
@@
The rise of deep learning (DL) models and the growing availability of extensive weather data sets are ushering in new set of tools and new opportunities for improved weather forecasting (Bouallegue et al. (2024)). Following a similar trend in other areas, several foundation weather forecasting models have recently appeared with striking capabilities (Pathak et al. (2022); Lam et al. (2022); Bi et al. (2023); Chen et al. (2023)). However precipitation remains poorly handled by these foundation AI models, and is often omitted from their evaluations. In some of these models, the approach adopted may not be well-suited for precipitation forecasting. For instance the dynamical system perspective adopted in GraphCast (Lam et al. (2022)) using a mesh-based graph convolutional neural networks is predicated on a smooth dynamics of the atmosphere, which typically does not apply to rainfall under the tropics.
In this work we tackle the tropical rainfall prediction problem at a smaller, region scale. Unlike foundation models, we focus solely on rainfall prediction, and we build models that are more modest in size and can be easily deployed with limited resources. Specifically, using the ERA5 and GPM-IMERG datasets we develop two deep learning (DL) models for 24h rainfall prediction over Ghana, with prediction lead-times of 12 hours and 30 hours respectively. Overall, we found that our 12h lead-time model matches, and in some accounts outperforms the 18h lead-time forecasts by the ECMWF (as available in the TIGGE dataset). Another important contribution of our work is that we develop a statistical method to probe the relative importance of the meteorological variables used in our model, leading to useful insights into the factors driving precipitation in the Ghana region.
@@
Beside foundation models, several other data-driven models have been developed for weather and precipitation prediction at a regional scale (Shi et al., 2015; Dueben and Bauer, 2018; Agrawal et al., 2019; Weyn et al., 2019; Ayzel et al., 2020; Ko et al., 2022; Oh et al., 2023). Several works have also targeted specifically rainfall prediction in West africa using data-driven approaches (Vogel et al. (2020, 2021); Gebremichael et al. (2022); Ageet et al. (2023)), including (Walz et al., 2024a) that uses deep learning techniques and is the closest to our work. Compared to (Walz et al., 2024a), the prediction problem that we consider in this work is more challenging. Indeed, our work focuses on Ghana at a 64-by-64 spatial resolution, compared to a 19-by-61 resolution in (Walz et al., 2024a) for an area 25 times bigger than Ghana. The challenge in rainfall prediction increases significantly as the spatial resolution is refined. Furthermore, we consider predictions with longer lead-times (12h and 30h compared to 6h in (Walz et al., 2024a)). Another important contribution of our work is a novel and widely applicable methodology that has allowed us to assess the relative importance of the meteorological variables used.
@@
The rest of the manuscript is structured as follows. Section 2 presents the general modeling approach and the datasets employed, Section 3 discusses our methodology, specifically the U-Net architecture employed. Section 4 describes the evaluation setup, including a novel approach for probing the importance of the input variables. In Section 5 we present and discuss our finding. Finally Section 6 contains some concluding thoughts.
@@ 2 General approach and datasets
The current study focuses on Ghana, a country in West Africa spanning latitudes $4.3^{\circ}$ North to $11.6^{\circ}$ North and longitudes $3.8^{\circ}$ West to $1.8^{\circ}$ East, as depicted on Figure 1. By area size, the country is smaller but comparable to the United Kingdom, and to the state of Michigan in the United States. The climate of Ghana is tropical, with distinct rainy and dry seasons, with the rainy season usually lasting from April to November, followed by a dry season (Manzanas et al., 2014). The rainy season varies with the latitudes: the south typically experiences two rainy seasons (a big rainy season from April to July, and a small one from September to November), whereas the north typically experiences one rainy season from April to October).
Figure 1: Ghana: The area of Interest within West Africa.
For this study daily rainfall measurements and meteorological variables over Ghana were collected from the GPM-IMERG data product developed by the National Aeronautics and Space Administration (NASA), and from the ERA5 data product of the European Center for Mid-range Weather Forecasting (ECMWF). We describe these data products below. Here we give a general overview of our approach. Specifically, our study time-window is June 1st, 2000 to September 30th, 2021, and we discretize the bounding box around Ghana into a $64\times 64$ image. For each date $t$ in the study window, we let ${\bf y}_{t}=\{y_{i,t},\;i\in 64\times 64\}$ be the image of rainfall observations over the 24h of date $t$, obtained from the GPM-IMERG data product. Actually, we follow the convention of GPM-IMERG that defined the 24h rainfall period as 6AM-6AM UTC. Similarly, for each date $t$ in the study window, and for a prediction with lead-time $h$, we collected $K=57$ environmental and meteorological variables over Ghana from the ERA5 data product at time $t-h$, as described below. We denote these variables on pixel $i$ by ${\bf x}_{i,t-h}\in\mathbb{R}^{K}$, and we set ${\bf x}_{t-h}=\{{\bf x}_{i,t-h},\;i\in 64\times 64\}$ for the 3D images of meteorological predictor variables.
@@ 2.1 GPM-IMERG precipitation data
The Global Precipitation Measurement (GPM) mission is an international satellite mission launched jointly by NASA and the Japanese JAXA in 2014 for precipitation measurements at the scale of the planet (Hou et al. (2014)). The mission was built on the successful Tropical Rainfall Measuring Mission (TRMM) satellites that operated from 1997 to 2015. The GPM mission carries the first space-borne Ku/Ka-band Dual-frequency Precipitation Radar (DPR) and a multi-channel GPM Microwave Imager (GMI). The Integrated Multi-satellite Retrievals for the Global Precipitation Measurement (IMERG) is a multi-satellite data processing algorithm developed by NASA (Huffman et al. (2020)). The algorithm is applied to the data collected by TRMM and GPM, and further leverages whatever constellation of satellites is available at a given time to produce a long time-series of rainfall measurement at the scale of the planet. The GPM-IMERG product is the current state-of-the-art satellite precipitation observation system, particularly under the tropics. The data is available at a 30-minute temporal resolution and 0.1 degree spatial resolution. For this study, we collected 24h accumulated (6AM-6AM UTC) GPM-IMERG data over the bounding box of Ghana at 0.1 degree spatial resolution, that we then regrid to a $64\times 64$ image. The image obtained for date $t$ is denoted ${\bf y}_{t}$. Although GPM-IMERG is known to have some bias, we will consider it as our precipitation ground truth.
@@ 2.2 ERA5 meteorological variables
The ECMWF Reanalysis version 5 (ERA5) is the 5th generation of the gridded reanalysis climate and weather dataset maintained by the European Center for Mid-range Weather Forecasting (ECMWF). See e.g. (Hersbach et al., 2020) for a description. The dataset is produced through reanalysis, that is through a data assimilation approach that combines numerical weather models with global climate observations. ERA5 data contains a large number of environmental and atmospheric variables, from 1940 onwards, at an hour temporal resolution, and $0.25^{\circ}$ spatial resolution. The ERA5 dataset is driving much of the recent machine learning effort for data-driven weather forecasting over the last several years (Keisler (2022); Lam et al. (2022); Chen et al. (2023)). We note that the dataset is not available in real-time. Its current latency is 5 days. This means that our model – like most other data-driven weather prediction models that rely on ERA5 data as inputs – is not currently operationally implementable.
Table 1 lists the 55 ERA5 variables used for this project. We compiled this list from the literature, notably Walz et al. (2024a), as well as from a basic correlation analysis. We group the variables following Walz et al. (2024a). As noted in Walz et al. (2024a), all these variables have been reported in the literature as playing some role in rain formation processes. We retrieved these variables over the bounding box around the Ghana region: $(3.8^{\circ}W,1.8^{\circ}E)$, and $(11.6^{\circ}N,4.3^{\circ}S)$, and with a temporal resolution of six hours.
@@
Since Ghana has a seasonal rainfall pattern that also varies strongly with latitude, accounting for space-time variability in the model is crucial for accurate forecasts. Thus in addition to the 55 variables listed in Table 1, for each pixel $i$ and each date $t$, we also constructed two new variables
$\textsf{COS}_{i,t}=\cos\left(2\pi\frac{\mathsf{day}(t)}{365}\right)\times\mathsf{lat}(i),$ (1)
$\textsf{SIN}_{i,t}=\sin\left(2\pi\frac{\mathsf{day}(t)}{365}\right)\times\mathsf{lat}(i),$ (2)
where $\mathsf{day}(t)\in\{1,\ldots,365\}$ is the day of the year of date $t$, and $\mathsf{lat}(i)$ is the latitude of the center point of pixel $i$.
In summary, for each date $t$ in our study window (between June 1st, 2000 and September 30th, 2021), and for each pixel $i\in 64\times 64$, we collected ${\bf x}_{i,t-h}\in\mathbb{R}^{57}$, composed of the values of the 55 variables listed in Table 1, plus the two constructed variables $\mathsf{COS}$ and $\mathsf{SIN}$, $h$ hours before 6AM UTC of date $t$.
@@ 2.3 TIGGE forecast data
We compare the skills of our models with the predictions from the European Center for Mid-range Weather Forecasting (ECMWF) made available through the THORPEX Interactive Grand worldwide Ensemble (TIGGE) program (Swinbank et al. (2016)). The TIGGE database has established itself as a key window into the capability of state-of-the-art operational NWP models, and has proven very useful to the research community. For this project we use the Noon-forecasts of the ECMWF model, which corresponds to a 18h lead-time 24h rainfall forecast. The ECMWF forecast that would match our 12h lead-time is not available in TIGGE. We downloaded their total rainfall (the sum of large scale organized precipitation and convective precipitation), that we re-grid to $64\times 64$ images.
@@ 3 Methodology
Using the same U-Net architecture, we trained two models with different lead-times. The first model has a lead-time $h=12$ hour that we denote by $\mathsf{UNET}_{12}$ (this model uses as input, the state of the meteorological variable at 6PM, 12 hours before the 6AM-6AM rainfall window). The second model has a lead-time of $h=30$ hour, and is denoted $\mathsf{UNET}_{30}$ (based on meteorological variables at Midnight, 30 hours before the 6AM-6AM rainfall window).
Given the training data $\mathcal{D}^{\prime}$, we train the U-Net models (3) by minimizing the L1 loss
$\textsf{L}(W;\mathcal{D})=\sum_{t\in\mathcal{D}^{\prime}}\;\left\|\textbf{y}_{t}-\mathcal{F}_{W}\left({\bf x}_{t-h}\right)\right\|_{1},$ (4)
where for an image ${\bf a}$, $\|{\bf a}\|_{1}$ denotes the sum of the absolute values of the pixel values of ${\bf a}$. The L1 loss is used for added robustness to outliers. We solve this minimization problem using the Adam optimizer (Kingma and Ba (2015)) with a learning rate of $1e^{-4}$ and a weight decay of $1e^{-4}$. During training, a batch size of $256$ is used (based on the availability of the GPU memory). We adjust the learning rate dynamically during training, using the ReduceLROnPlateau method Krizhevsky et al. (2012). This scheduler monitors the loss and adjusts the learning rate when a plateau in performance is detected. Specifically, the learning rate is reduced by a factor of $0.1$ if no improvement is observed in the loss for a specified patience period of $100$ epochs.
@@
We evaluate the performance of the model by comparing $\hat{\bf y}_{t^{\prime}}$ and ${\bf y}_{t^{\prime}}$ over the test set $\mathcal{D}^{{}^{\prime\prime}}$, as we describe below.
3.1 Combining $\mathsf{UNET}_{12}$ and ECMWF forecasts
In addition of the two models $\mathsf{UNET}_{12}$ and $\mathsf{UNET}_{30}$ we also explore a prediction model that combines $\mathsf{UNET}_{12}$ with the 18h lead-time forecast of the ECMWF obtained from the Tigge database. This averaging approach takes stock in the distinctive features of both methodologies, potentially yielding a more accurate forecast. To accomplish this, and after some trial-and-errors, a weighted average of the ECMWF forecast (with weight $0.1$), and the $\mathsf{UNET}_{12}$ prediction (with weight $0.9$) is proposed. The resulting ensemble weighted average forecast will be denoted below by $\mathsf{Ens}$.

### v3

@@ Abstract
With a warming planet, tropical regions are expected to experience the brunt of climate change, with more intense and more volatile rainfall events. Currently, state-of-the-art numerical weather prediction (NWP) models are known to struggle to produce skillful rainfall forecasts in tropical regions of Africa. There is thus a pressing need for improved rainfall forecasting in these regions. Over the last decade or so, the increased availability of large-scale meteorological datasets and the development of powerful machine learning models have opened up new opportunities for data-driven weather forecasting. Focusing on Ghana in this study, we use these tools to develop two U-Net convolutional neural network (CNN) models to predict 24h rainfall at 12h and 18h lead-time. The models were trained using data from the ERA5 reanalysis dataset and the GPM-IMERG dataset. A special attention was paid to interpretability. We developed a novel statistical methodology that allowed us to probe the importance of the meteorological variables input in our model, offering useful insights into the factors that drive precipitation in the Ghana region. Empirically, we found that our 12-hour lead-time model consistently delivers the strongest performance, and both our 12-hour and 18-hour lead-time models outperform the climatology forecast in every aspect. Additionally, their performance matches and sometimes surpasses the 18-hour lead-time forecasts produced by the ECMWF ensemble forecast (as available in the TIGGE dataset).
@@ 1 Introduction
The precipitation of liquid water from Earth’s atmosphere to the surface is one of the most crucial atmospheric processes shaping life on Earth (Box and Box, 2015). Lack of rainfall leads to drought, and too much rainfall leads to flooding, two extreme weather events that often lead to natural disasters. The spatio-temporal distribution and intensity of precipitation events also hold significant implications for the smooth running of human societies, from day-to-day events, to large-scale urban and agricultural plans. As such, precipitation prediction is an essential tool for planning and decision making. However, rain and other forms of precipitation are notoriously difficult to predict (Tapiador et al., 2019). Unlike most atmospheric variables, rainfall amounts are spatio-temporally non-smooth, and this poses a significant challenge to modeling, with many models often based (implicitly or explicitly) on smoothness assumptions.
@@
The situation is particularly problematic under the tropics where rainfall is dominated by multiscale convective flows that are currently challenging to represent in numerical weather prediction (NWP) models (Becker et al., 2021). In tropical West Africa, this issue is compounded with the lack of quality measurements. As a consequence, state-of-the-art NWP models such as the European Center for Medium-range Weather Forecasts (ECMWF) models are known to struggle to produce skillful forecasts in these regions (Vogel et al., 2020; Kniffka et al., 2020; Rojas-Campos et al., 2023). At the same time, regions of sub-Saharan Africa are projected to experience the brunt of climate change, with more intense and more volatile precipitation as Earth’s atmosphere warms (WMO, 2024). For instance, flooding from the 2024 rainy season in West and Central Africa has claimed the lives of more than a thousand, and displaced close to a million people [^1] https://www.nytimes.com/2024/09/15/world/africa/floods-africa.html?smid=url-share.
@@
The rise of deep learning (DL) models and the growing availability of extensive weather data sets are ushering in a new set of tools and new opportunities for improved weather forecasting (Bouallegue et al., 2024). Following a similar trend in other areas, several foundation weather forecasting models have recently appeared with striking capabilities (Kurth et al., 2023; Lam et al., 2023; Bi et al., 2023; Chen et al., 2025). However, precipitation remains poorly handled by these foundation AI models and is often omitted from their evaluations. In some of these models, the approach adopted may not be well-suited for precipitation forecasting. For instance, the dynamical system perspective adopted in $\rm{GraphCast}$ uses a mesh-based graph convolutional neural network and is predicated on the smooth dynamics of the atmosphere (Lam et al., 2023), which typically does not apply to rainfall in the tropics.
@@
Beside foundation models, several other data-driven models have been developed for weather and precipitation prediction at a regional scale (Shi et al., 2015; Dueben and Bauer, 2018; Agrawal et al., 2019; Weyn et al., 2019; Ayzel et al., 2020; Ko et al., 2022; Oh et al., 2023). Several works have also targeted specifically rainfall prediction in Sub-Saharan Africa using data-driven approaches (Vogel et al., 2020, 2021; Gebremichael et al., 2022; Ageet et al., 2023), including Walz et al. (2024a), which uses deep learning techniques and is the closest to our work. Compared to Walz et al. (2024a), the prediction problem that we consider in this work is more challenging as we discuss below.
In this work, we tackle the tropical rainfall prediction problem at a smaller, regional scale. Unlike foundation models, we focus solely on rainfall prediction, and we build models that are more modest in size and can be easily deployed with limited resources. Specifically, using the 5th-generation Reanalysis (ERA5) of the ECMWF and the 6th-generation Integrated Multi-satellitE Retrievals for Global Precipitation Measurement (GPM-IMERG) datasets, we develop two deep learning (DL) models for 24h rainfall prediction over Ghana, with prediction lead-times of 12 hours and 18 hours, respectively. Overall, we found that our 12-hour lead-time model consistently delivers the strongest performance, and both our 12-hour and 18-hour lead-time models outperform the climatology forecast in every aspect. Additionally, we observed that using a hybrid approach that combines our data-driven model with classical NWP techniques further enhances forecast accuracy.
Our work focuses on an area $7.3^{\circ}$ by $5.6^{\circ}$ large, centered around Ghana at a 64-by-64 spatial resolution, compared to a 19-by-61 resolution in Walz et al. (2024a) for an area 25 times bigger. The challenge in rainfall prediction increases significantly as the spatial resolution is refined. Furthermore, we consider predictions with longer lead-times (12h and 18h compared to 6h in Walz et al. (2024a)). Another important contribution of our work is that we develop a statistical method to probe the predictive importance of the meteorological variables used in our model, leading to useful insights into the factors driving precipitation in the Ghana region.
@@
The rest of the manuscript is structured as follows. Section 2 presents the general modeling approach and the datasets employed, and Section 3 discusses our methodology, specifically the U-Net architecture employed. Section 4 describes the evaluation setup, including a novel approach for probing the importance of the input variables. In Section 5 we present and discuss our findings. Finally, Section 6 contains some concluding thoughts.
@@ 2 General approach and datasets
The current study focuses on an area centered around Ghana in West Africa, and spanning the latitudes $4.3^{\circ}$ North to $11.6^{\circ}$ North and the longitudes $3.8^{\circ}$ West to $1.8^{\circ}$ East, as depicted on Figure 1. The climate of Ghana is tropical, with distinct rainy and dry seasons, with the rainy season usually lasting from April to November, followed by a dry season (Manzanas et al., 2014). The rainy season varies with the latitudes: the south typically experiences two rainy seasons – a long rainy season from April to July, and a short one from September to November - whereas the north typically experiences one rainy season from April to October.
Figure 1: Ghana: The area of interest within Africa.
Source: Google. (n.d.). Partial map of Africa. Retrieved April 21st, 2025, using the Google map platform.
[…]
@@ 2.1 GPM-IMERG precipitation data
The Global Precipitation Measurement (GPM) mission is an international satellite mission launched jointly by NASA and the Japanese JAXA in 2014 for precipitation measurements at the scale of the planet (Hou et al., 2014). The mission was built on the successful Tropical Rainfall Measuring Mission (TRMM) satellites that operated from 1997 to 2015. The GPM mission carries the first space-borne Ku/Ka-band Dual-frequency Precipitation Radar (DPR) and a multi-channel GPM Microwave Imager (GMI). The Integrated Multi-satellite Retrievals for the Global Precipitation Measurement (IMERG) is a multi-satellite data processing algorithm developed by NASA (Huffman et al., 2020). The algorithm is applied to the data collected by TRMM and GPM, and further leverages whatever constellation of satellites is available at a given time to produce a long time-series of rainfall measurements at the scale of the planet (Huffman et al., 2020). The GPM-IMERG product is the current state-of-the-art satellite precipitation observation system, particularly under the tropics. The data is available at a 30-minute temporal resolution and 0.1-degree spatial resolution. For this study, we downloaded the 30-minute rainfall data and derived 24h accumulated (6AM-6AM UTC) GPM-IMERG version 6 (V6) data over the bounding box of Ghana at 0.1 degree spatial resolution, which we then regrid (cubic interpolation) to a $64\times 64$ image. The image obtained for date $t$ is the previously introduced ${\bf y}_{t}$. Although GPM-IMERG is known to have some bias (Pradhan et al., 2022), we will consider it as our precipitation ground truth.
@@ 2.2 ERA5 meteorological variables
The ECMWF Reanalysis version 5 (ERA5) is the 5th generation of the gridded reanalysis climate and weather dataset maintained by ECMWF (Hersbach et al., 2020). The dataset is produced through reanalysis, that is, through a data assimilation approach that combines numerical weather models with global climate observations at hourly temporal resolution. ERA5 data contains a large number of environmental and atmospheric variables, from 1940 onwards, at an hour temporal resolution, and $0.25^{\circ}$ spatial resolution. The ERA5 dataset is driving much of the recent machine learning effort for data-driven weather forecasting (Keisler, 2022; Lam et al., 2023; Chen et al., 2025). We would like to note that the dataset is not available in real-time, as it currently has a latency of 5 days. As a result, our model, similar to most other data-driven weather prediction models that depend on ERA5 data inputs, is not yet operationally implementable.
Table 1 lists the 55 ERA5 variables used for this project. We compiled this list from the literature, notably Walz et al. (2024a), as well as from a basic correlation analysis. We group the variables following Walz et al. (2024a). As noted in Walz et al. (2024a), all these variables have been reported in the literature as playing some role in rain formation processes. We retrieved these variables over the area of interest with a temporal resolution of six hours.
@@
Since Ghana has a seasonal rainfall pattern that also varies strongly with latitude, accounting for space-time variability in the model is crucial for accurate forecasts. Thus in addition to the 55 variables listed in Table 1, for each pixel $i$ and each date $t$, we also constructed two new variables:
$\rm{COS}_{i,t}=\cos\left(2\pi\frac{\rm{day}(t)}{365}\right)\times\rm{lat}(i),$ (1)
$\rm{SIN}_{i,t}=\sin\left(2\pi\frac{\rm{day}(t)}{365}\right)\times\rm{lat}(i),$ (2)
where $\rm{day}(t)\in\{1,\ldots,365\}$ is the day of the year of date $t$, and $\rm{lat}(i)$ is the latitude of the center point of pixel $i$.
In summary, for each date $t$ in our study window (between June 1st, 2000 and September 30th, 2021), and for each pixel $i\in 1\dv 64\times 1\dv 64$, we collected ${\bf x}_{i,t-h}\in\mathbb{R}^{57}$, composed of the values of the 55 variables listed in Table 1, plus the two constructed variables $\rm{COS}$ and $\rm{SIN}$, $h$ hours before 6AM UTC of date $t$. It is worth mentioning that precipitation was not included as an input variable in the model. This allows our model to learn meaningful relationships between the available meteorological variables and rainfall, without reliance on a mechanistic auto-regressive model.
@@ 2.3 TIGGE forecast data
We compare the skills of our models with the predictions from ECMWF, made available through the THORPEX Interactive Grand Global Ensemble (TIGGE) program (Swinbank et al., 2016). The TIGGE database has established itself as a key resource for evaluating the capabilities of state-of-the-art operational NWP models and has proven invaluable to the research community. For this project, we use the Noon forecasts from the ECMWF ensemble (50 perturbed members + 1 control member) available through TIGGE. The ECMWF noon-forecast predicts the state of the atmosphere every 6h starting from Noon. From these forecasts, we derived the predicted 24h accumulation of rainfall from 6AM (next day) to 6AM the day after next, corresponding to an 18h lead-time prediction. We focus on 6AM-6AM to match the GPM-IMERG precipitation window. We downloaded the total rainfall (the sum of large-scale organized precipitation and convective precipitation) from the ECMWF ensemble forecasts and re-gridded it to $64\times 64$ images for further analysis.
@@ 3 Methodology
The U-Net architecture is a CNN-based deep neural network (DNN) architecture specifically designed for image tasks. Specifically, we utilize a 2D U-Net model with three downsampling blocks, each containing two convolutional layers, followed by three upsampling blocks with concatenation skip connections (see Figure 2). In contrast to more computationally expensive architectures like graph neural networks (GNNs) and Transformers, U-Net models offer a balance between accuracy and efficiency, and have been used for weather data in several prior works (Lam et al., 2023; Keisler, 2022; Walz et al., 2024a). The model leverages an encoder-decoder structure with concatenation skip connections to achieve high segmentation accuracy while maintaining computational efficiency. The encoder layers map the input data to a more abstract/latent space. After further abstraction through the convolutions in the middle layers, the decoder layers map the data back to a gridded image. The skip connections concatenate same-level encoding and decoding layers, which is believed to improve performance for image-to-image problems (Ronneberger et al., 2015), although a rigorous understanding remains an open problem.
@@
Using the same U-Net architecture, we trained two models with different lead-times. The first model has a lead-time $h=12$ hour that we denote by UNET12 (this model uses as input, the state of the meteorological variable at 6 PM, 12 hours before the 6AM-6AM rainfall window). The second model has a lead-time of $h=18$ hours, and is denoted UNET18 (based on meteorological variables at Noon, 18 hours before the 6AM-6AM rainfall window).
Given the training data $\mathcal{D}^{\prime}$, we train the U-Net models (3) by minimizing the L1 loss in (4) with respect to $W$:
$\rm{L}(W;\mathcal{D}^{\prime})=\sum_{t\in\mathcal{D}^{\prime}}\;\left\|\textbf{y}_{t}-\mathcal{F}_{W}\left({\bf x}_{t-h}\right)\right\|_{1},$ (4)
where for an image ${\bf a}$, $\|{\bf a}\|_{1}$ denotes the sum of the absolute values of the pixel values of ${\bf a}$. The L1 loss is used for added robustness to outliers. We solve this minimization problem using the Adam optimizer (Kingma and Ba, 2015) with a learning rate of $10^{-4}$ and a weight decay of $10^{-4}$. During training, a batch size of $256$ is used (based on the availability of the GPU memory). We adjust the learning rate dynamically during training, using the ReduceLROnPlateau method Krizhevsky et al. (2012). This scheduler monitors the loss and adjusts the learning rate when a plateau in performance is detected. Specifically, the learning rate is reduced by a factor of $0.1$ if no improvement is observed in the loss for a specified patience period of $100$ epochs.
@@
For each test date $t^{\prime}\in\mathcal{D}^{{}^{\prime\prime}}$, the prediction $\hat{\bf y}_{t^{\prime}}=\{\hat{y}_{i,t^{\prime}},\;i\in 1\dv 64\times 1\dv 64\}$ is a $64\times 64$ image, where $\hat{y}_{i,t}$ is the prediction at pixel $i$. We evaluate the performance of the model by comparing $\hat{\bf y}_{t^{\prime}}$ and ${\bf y}_{t^{\prime}}$ over the test set $\mathcal{D}^{{}^{\prime\prime}}$, as we describe in the next section.
3.1 Uncertainty quantification using EasyUQ
For uncertainty quantification, and for a meaningful comparison with the TIGGE ensemble predictions, we use the EasyUQ method developed by Walz et al. (2024b) to estimate the cumulative distribution function (cdf) of the conditional distribution of the ground truth $y_{i,t}$ given $\hat{y}_{i,t}$. This provides a measure of uncertainty on the prediction $\hat{y}_{i,t}$.
These estimated CDFs are obtained by solving an isotonic regression problem on the training dataset, with interpolation adjustments on the test dataset. We give a brief description of the method, and refer the reader to Henzi et al. (2021); Walz et al. (2024b) for more details, and for software. Fix a pixel $i$, and let $\{(y_{i,t},\hat{y}_{i,t}),\;t\in\mathcal{D}^{{}^{\prime}}\}$ be the ground truths and prediction values in the training set at pixel $i$. Let $F_{\hat{y}_{i,t}}$ denote the conditional cdf of $y_{i,t}$ given $\hat{y}_{i,t}$. EasyUQ estimates the map $\{\hat{y}_{i,t},\;t\in\mathcal{D}^{{}^{\prime}}\}\mapsto\{F_{\hat{y}_{i,t}},\;t\in\mathcal{D}^{{}^{\prime}}\}$ (a distributional regression problem) by minimizing the loss
$\min_{\{F_{i,t},\;t\in\mathcal{D}^{{}^{\prime}}\}}\;\sum_{t\in\mathcal{D}^{{}^{\prime}}}{\rm CRPS}(F_{i,t},y_{i,t}),\;\;\;\;\mbox{ where }\;\;\;\;{\rm CRPS}(F,y)=\int_{-\infty}^{+\infty}\left(F(u)-\textbf{1}_{\{y\leq u\}}\right)^{2}\mathrm{d}u,$
is the continuous ranked proper scoring (CRPS) of cdf $F$ and target $y$ (see Gneiting and Raftery (2007) for more details on the CRPS). The minimization problem described above is a minimization over cdfs $\{F_{i,t},\;t\in\mathcal{D}^{{}^{\prime}}\}$. Under a monotonicity assumption, it was shown by Henzi et al. (2021) that this minimization problem has a unique solution $\{\hat{F}_{i,t},\;t\in\mathcal{D}^{{}^{\prime}}\}$ that can be easily computed. $\hat{F}_{i,t}$ is the EasyUQ estimate of $F_{\hat{y}_{i,t}}$. When apply on the sample $t^{\prime}$ in the test set $\mathcal{D}^{{}^{\prime\prime}}$, $F_{\hat{y}_{i,t^{\prime}}}$ is estimated by interpolation from the estimates $\{\hat{F}_{i,t},\;t\in\mathcal{D}^{{}^{\prime}}\}$ in the training set.
3.2 Combining UNET18 and ECMWF ensemble forecasts
In addition to the two models UNET12 and UNET18 we also explore a prediction model that combines UNET18 with the 18h lead-time forecast of the ensemble ECMWF (referred as NWP forecast) obtained from the TIGGE database. This hybrid approach takes stock of the distinctive features of both methodologies, potentially yielding a more accurate forecast. For $m\in\{{\rm UNET}_{18},\;{\rm NWP}\}$, we write $\hat{y}_{i,t}(m)$ to denote the prediction of method $m$ for time $t$ and pixel $i$. We define the hybrid prediction as
$\hat{y}_{i,t}(\mathrm{HYB})=(1-\beta)\cdot\hat{y}_{i,t}(\mathrm{UNET}_{18})+\beta\cdot\hat{y}_{i,t}(\mathrm{NWP}),$ (5)
where the $\beta$ is estimated via linear regression over the test set by regressing the true residual $y_{i,t}-\hat{y}_{i,t}(\mathrm{UNET}_{18})$ on the difference $\hat{y}_{i,t}(\mathrm{NWP})-\hat{y}_{i,t}(\mathrm{UNET}_{18})$, without an intercept. We obtain $\beta=0.546$, indicating that the hybrid is almost an equal-weight average of both models, with slightly more weight given to the NWP forecast.
We obtain an ensemble of hybrid predictions by applying Equation (5) to each of the 51 members of the NWP ensemble. Note that the EasyUQ method of Walz et al. (2024b) cannot be applied to the hybrid prediction here, since we do not have the TIGGE predictions on the training dataset.